Ramon Perez worked 40 hours at $13.00 an hour and five hours of overtime at time-and-a-half. What was his pay for the week?
Answer by JulietG(1812)

Ramon earned:
40 * 13 = 520
5 * 13(1.5) =
5 * 19.50 = 97.50
520 + 97.50 = $617.50
.
Of course, Ramon doesn't actually get to take home $617.50. A substantial portion of his check will be withheld for taxes so that he can pay for Julio to sit on his behind all day and collect welfare.